OpenStudy (anonymous):

can u explain how....-3×(-2)=6

OpenStudy (anonymous):

Explain please.......

OpenStudy (uri):

Is that a mutiply sign or it's a 3x?

OpenStudy (anonymous):

Multiply

OpenStudy (anonymous):

-3×(-2)=6..........how

OpenStudy (uri):

-3×(-2)=6 6=6

OpenStudy (uri):

Move the other six at right side 6-6 they both will be cancelled answer is 0

OpenStudy (anonymous):

Please ....question is explain that -3 × (-2) = 6, how step by step please.

OpenStudy (uri):

-3x-2 is 6.

OpenStudy (anonymous):

Is it not like this that when -2, comes out of (-2), it becomes +2....

OpenStudy (uri):

No...when there are two - like you can see -3 and -2 will be - and - + so 3x2 =6

OpenStudy (anonymous):

So what is the meaning of ( )

OpenStudy (uri):

Nothing really...it's just that when we have a postive +2 we can write like 2 when we have a -2 we have to write like -2 so that's how.

OpenStudy (anonymous):

So the answe is 0......

OpenStudy (uri):

Yes because 6=6 we take the other 6 to right side sign changes and it becomes negative 6 so 6-6 They both will be cancelled...so undefined..

OpenStudy (anonymous):

Thanks a lot .......sir/maadam

OpenStudy (uri):

Madam* :P

OpenStudy (uri):

Call me Uri.

OpenStudy (anonymous):

Thank u uri.....

OpenStudy (anonymous):

Bye.......ciao....